OpenStudy (russianboss1):

Medal and Fan! A quadratic equation has a discriminant of -1. Which answer describes the number and type of solutions of the equation? a)1 real and 1 complex solution b)1 real solution c)2 complex solutions d)2 real solutions

OpenStudy (welshfella):

Hint:- the square root of a negative number is not real

OpenStudy (welshfella):

Consider the quadratic formula x = [-b +/- sqrt(b^2 - 4ac)] / 2a (the discriminant is b^2 - 4ac)

OpenStudy (russianboss1):

I still don't get it

OpenStudy (welshfella):

b^2 - 4ac is = negative 1 so square root of this is imaginary Do you know what a complex number is?

OpenStudy (welshfella):

Its a number made up of a real part and an imaginary part - an example is 3 + 2i where i is the symbol for sqrt(-1)

OpenStudy (gottennis121):

@Englishguy @ransondyara471 @imqwerty

OpenStudy (gottennis121):

@pooja195 @Mckeanzie @-SpencerBraz- @mathmale

OpenStudy (welshfella):

let's substitute for the discriminant in the formula x = - b +/- sqrt(-1) / 2a the +/- indicates there are 2 solutions

OpenStudy (welshfella):

so x = ( -b + sqrt(-1)) / 2a or (-b - sqrt(-1)) / 2a sqrt(-1) = i so we have (-b + i ) / 2a and (- b - i)/ 2a = (-b/2a) + (i/2a) and (-b/2a) - (i/2a) these are 2 complex numbers -b/2a is the real part and -/2a is the imaginary part